OpenStudy (anonymous):

Differentiate y=c cos t + t^2 sin t

OpenStudy (anonymous):

@zepdrix

OpenStudy (anonymous):

help please and explain?

OpenStudy (anonymous):

cos t = - sin t sin t = cos t

OpenStudy (anonymous):

@dan815

OpenStudy (dan815):

Product rule of differentiation for the 2nd term

OpenStudy (dan815):

d/dt[t^2 sint] = f(t) * g(t) = f'(t) g(t) + f(t)*g'(t) = 2t sint + t^2*cost

OpenStudy (dan815):

d/dt [c cos t + t^2 sint] = -c sint + 2t sint + t^2*cost